# KANAZAWA-24-05

### v1

@@ 1 Introduction
Why new neutral leptons? To generate neutrino masses and lepton mixing as observed in neutrino oscillation phenomena, one needs to go beyond the SM. Many options are being explored as extensions of the Higgs and/or gauge sectors, mostly with new fields in the particle content. In particular, heavy neutral fermions, such as right-handed neutrinos ($\nu_{R}$), are often present as building blocks of several mechanisms for neutrino mass and lepton mixing generation. For example, in the case of the type-I seesaw mechanism [9, 10, 11, 12, 13], at least two $\nu_{R}$ are required to accommodate neutrino data and, in order to accommodate sub-eV neutrino masses (the most recent direct neutrino-mass measurement by KATRIN experiment [14] gives the upper limit of $\sim 0.8$ eV), the type-I seesaw mechanism requires either a high lepton number violating (LNV) scale, or very small Yukawa couplings if the the mechanism is realised at low scale, and is thus difficult to probe experimentally. An alternative to have a low-scale seesaw realisation with large Yukawa couplings is to consider additional sterile fermions and an approximate lepton number symmetry. This is the case of, for instance, the inverse seesaw (ISS) [15, 16, 17, 18, 19] or the linear seesaw (LSS) [20, 21] realisations, where a $B-L$ global symmetry is used to protect the active neutrino masses, linking their smallness to small parameters that quantify the breaking of the $B-L$ symmetry. In these mechanisms, where other additional sterile - from the SM gauge point of view - leptons ($\nu_{S}$) are considered, it is possible to obtain large neutrino Yukawa couplings with a comparatively small seesaw scale, potentially within the reach of colliders and also with potential impact on many observables like charged lepton EDMs and charged lepton flavour/number violating processes.
@@
In fact, this exact cancellation should also occur in the case of models with pseudo-Dirac sterile fermions since its origin is by virtue of the Ward-Takahashi identity in quantum electrodynamics. In this case, the contribution from the other Feynman diagrams are thus anticipated to make the dominant contribution compared to that of the rainbow-type diagrams. The aim of the present work is therefore to revisit the previous study of Ref. [23], taking into account the diagrams that had been set aside in the belief, prior to recent developments, that only rainbow-type Feynman diagrams made the dominant contribution to charged lepton EDMs, to update our calculation and to derive a complete analytical expression for them in the case of the inverse seesaw mechanism behind active light neutrino masses and lepton mixing generation.
@@
Our calculation is performed in a non-linear Feynman-’t Hooft gauge instead of the usual gauge fixing so that the number of the contributing Feynman diagrams is considerably reduced. After having derived the complete analytical expressions for the charged lepton EDMs, we conduct a thorough numerical analysis evaluating the electron EDM, complying with all relevant experimental bounds such as neutrino oscillation data, bounds on charged lepton flavour violating processes, sterile neutrino direct searches, meson decays, sterile neutrino decays, and cosmological and astrophysical observations. For this, we have chosen the case where the most minimal ISS mechanism is embedded into the SM, that is the (2,2) ISS [29], where only two pseudo-Dirac neutrinos contribute in the loops for the charged lepton EDMs. We note that this study is general to all seesaw models leading to pseudo-Dirac neutrino pairs in the heavy spectrum. We find that the magnitude of the predicted electron EDM is $\mathcal{O}(10^{-36})~{}e\hskip 1.42271pt\mathrm{cm}$ at most, which is smaller than the previous obtained result [23]. This is because the loop function in this updated calculation is smaller than before. Nevertheless, once more sterile neutrinos are considered, there are more contribution to the electron EDM but the enhancement is only $N(N-1)/2$ times larger than the most minimal case we considered here ($N=2$), where $N$ is the number of pairs of pseudo-Dirac neutrinos. For example, one can have a factor enhancement of 3 for $N=3$ and 6 for $N=4$. This enhancement is not enough to reach the future experimental sensitivity. In order to have an enhancement of a few orders of magnitude, one need to add a large number of sterile fermions, provided their masses are above $200$ MeV not to be in conflict with the severe constraint form Big-Bang nucleosynthesis [30]. If, in addition, one considers the ingredient fields at the origin of the LNV source, there might be further two-loop Feynman diagrams, where the heavy neutrinos as well as new neutral scalar fields run in the loops, contributing to the EDMs of charged leptons, further enhancing the phenomenological consequences of models involving sterile neutral leptons, see for instance [31].
@@ 3 Experimental constraints
Various constraints such as collider searches, meson decays, sterile neutrino decays, cosmological and astrophysical constraints are summarized in, for instance, Refs. [40, 41]. These constraints are translated into the upper bounds for the sterile neutrino mixing matrix elements $|U_{\alpha i}|^{2}$ as a function of the sterile neutrino masses. The bounds in Ref. [40] are interpreted as a pair of nearly degenerate states for the pseudo-Dirac sterile fermions, namely $|U_{\alpha 4}|^{2}+|U_{\alpha 5}|^{2}$ and $|U_{\alpha 6}|^{2}+|U_{\alpha 7}|^{2}$ in our case.
@@ 4.2 Analytical derivation of charged lepton EDM formulae
In the (approximate) lepton number conserving case, like the models with pseudo-Dirac fermions, there are three topological diagrams relevant to the charged lepton EDMs as shown in Fig. 1, where a photon can attach to any charged particles. [^5] For lepton number violating models such as Type-I seesaw, there are additional diagrams contributing to charged lepton EDMs [22]. In fact and following Ref. [50], one can show that the contribution from the third diagram of Fig. 1, that we name “rainbow” diagram, is exactly zero by virtue of the Ward-Takahashi identity in quantum electrodynamics. Indeed, it has been recently shown that the contribution from such rainbow diagrams vanishes exactly in some cases [27, 26]. In this work, we first check that the cancellation also occurs and is exact in our model with approximate lepton number symmetry and then we proceed to calculate the contribution of the remaining diagrams. The easiest way to show this cancellation in our case could be to perform the calculation using the unitary gauge, otherwise one has to evaluate a large number of diagrams and show the exact cancellation between them. Nevertheless, and for the sake of completeness, we have checked that the exact cancellation occurs in the non-linear Feynman-’t Hooft gauge. However, the latter calculation is tedious and time-consuming, and not worth showing in this study. The relevant rainbow Feynman diagrams are listed in Fig. 2, and the exact cancellations occur in each pair of diagrams R1–R4.
@@
The $h$ and $Z$ mediating loop contributions, are given by
@@ 5 Numerical calculations
Fig. 6 shows the obtained CP phase factor $|J_{\alpha}^{D}|$ for the electron case ($\alpha=e$) satisfying all the relevant constraints as a function of the lightest sterile fermion mass $m_{4}~{}(\approx m_{5})$ for Normal (left) and Inverted (right) orderings. One can find that the CP phase factor for the Inverted ordering case can be about four times larger than the one in the Normal ordering case, and the maximal magnitude is $\mathcal{O}(10^{-8})$ around $m_{4}\approx 100~{}\mathrm{GeV}$. For sterile fermion masses heavier than $\mathcal{O}(1)~{}\mathrm{TeV}$, the charged lepton flavour violating processes, in particular $\mu\to e\overline{e}e$, strongly constrain the parameter space due to the non-decoupling effects [39]. For $m_{4}\lesssim 10~{}\mathrm{GeV}$, the parameter region is also strongly limited by the beam dump experiments [40, 52], and thus we do not consider this region in our calculations.
@@ 6 Conclusions
CP-violating phases in the lepton sector are important for generating the baryon asymmetry of the Universe via leptogenesis are linked to the mechanisms for generating small neutrino masses and their mixing as observed in neutrino oscillation experiments. In this work, we considered the charged lepton EDMs at the two-loop level as a CP-violating observable in neutrino models based on approximate lepton number symmetry and thus presenting pseudo-Dirac sterile fermions in the heavy neutrino spectrum. We found that a number of diagrams relevant to the charged lepton EDMs exactly cancel (rainbow Feynman diagrams), and we computed all contributing two-loops diagrams in the non-linear Feynman-’t Hooft gauge to derive analytic formulae for the charged lepton electric dipole moments.
@@
In addition, we numerically evaluated the predicted charged lepton EDMs taking into account the current experimental constraints such as the neutrino oscillation data, bounds on the charged lepton flavour violating processes $\ell_{\alpha}\to\ell_{\beta}\gamma$ and $\mu\to e\overline{e}e$, sterile neutrino direct searches, meson decays, sterile neutrino decays, cosmological and astrophysical observations. As a result, we found that the predicted electron EDM is at most $\sim 5\times 10^{-36}~{}e\hskip 1.42271pt\mathrm{cm}$, which is too small to be explored by near future experiments. For muon and tau, the predicted EDMs are also much smaller than the experimental reaches. On the other hand, there are novel proposals aiming at measuring the electron EDM as small as $|d_{e}|<\mathcal{O}(10^{-35}-10^{-37})~{}e\hskip 1.42271pt\mathrm{cm}$ [5].
@@
This result, as is the analytical formula we derive, is generic to any fermionic seesaw based on approximate lepton number symmetry and leading to two pseudo-Dirac pairs in the heavy neutral spectrum, like, for instance, the most minimal inverse seesaw or linear seesaw mechanisms. Nevertheless, if we consider a complete model in which the seesaw mechanism is embedded, there are other two-loop contributions that will enhance these predictions. For instance, if more sterile neutrinos are considered, there are more contribution to the charged lepton EDMs but the enhancement is only $N(N-1)/2$ times larger than the most minimal case (with $N=2$), where $N$ is the number of pseudo-Dirac pairs, meaning a factor enhancement of 3 for $N=3$ and 6 for $N=4$, which is not enough to reach the future experimental sensitivity. In order to have an enhancement of a few orders of magnitude, one needs to add a large number of sterile fermions, provided their masses are above $\sim 200$ MeV not to be in conflict with the severe constraint form Big-Bang nucleosynthesis [30]. If, in addition, one considers a more complete model with the other ingredient fields at the origin of the lepton number violating source, there might be further two-loop Feynman diagrams, where the heavy neutrinos, as well as, new neutral and charged scalar fields run in the loops, contributing to the charged lepton EDMs and further enhancing the predicted values.
@@ Appendix A Non-linear Feynman-’t Hooft gauge
with $\nu_{\alpha}=U_{\alpha i}\nu_{i}$, $\theta_{W}$ is the Weinberg angle, $e=g\sin\theta_{W}$, $m_{W}=g\langle H^{0}\rangle/\sqrt{2}=m_{Z}\cos\theta_{W}$ and $\epsilon_{123}=+1$, the Lagrangian terms are given as follows

### v2

@@ 1 Introduction
Why new neutral leptons? To generate neutrino masses and lepton mixing as observed in neutrino oscillation phenomena, one needs to go beyond the SM. Many options are being explored as extensions of the Higgs and/or gauge sectors, mostly with new fields in the particle content. In particular, heavy neutral fermions, such as right-handed neutrinos ($\nu_{R}$), are often present as building blocks of several mechanisms for neutrino mass and lepton mixing generation. For example, in the case of the type-I seesaw mechanism [9, 10, 11, 12, 13], at least two $\nu_{R}$ are required to accommodate neutrino data and, in order to accommodate sub-eV neutrino masses (the most recent direct neutrino-mass measurement by KATRIN experiment [14] gives the upper limit of $\sim 0.8$ eV), the type-I seesaw mechanism requires either a high lepton number violating (LNV) scale, or very small Yukawa couplings if the mechanism is realised at low scale, and is thus difficult to probe experimentally. An alternative to have a low-scale seesaw realisation with large Yukawa couplings is to consider additional sterile fermions and an approximate lepton number symmetry. This is the case of, for instance, the inverse seesaw (ISS) [15, 16, 17, 18, 19] or the linear seesaw (LSS) [20, 21] realisations, where a $B-L$ global symmetry is used to protect the active neutrino masses, linking their smallness to small parameters that quantify the breaking of the $B-L$ symmetry. In these mechanisms, where other additional sterile - from the SM gauge point of view - leptons ($\nu_{S}$) are considered, it is possible to obtain large neutrino Yukawa couplings with a comparatively small seesaw scale, potentially within the reach of colliders and also with potential impact on many observables like charged lepton EDMs and charged lepton flavour/number violating processes.
@@
In fact, this exact cancellation should also occur in the case of models with pseudo-Dirac sterile fermions since its origin is by virtue of the Ward-Takahashi identity in quantum electrodynamics. This cancellation has been shown as below [26] . In general, one can write down the relevant diagrams contributing to the EDMs with the fermion self-energy $\Sigma(p/\hskip 0.85355pt)$ in the loop and the vertex correction $\Gamma^{\mu}(p,q)$. The transverse part of the vertex correction can be simplified with the Ward-Takahashi identity. In addition, one can also write down the longitudinal part of the vertex correction as a linear combination of possible vectors. Then, only the part giving a non-zero contribution to the EDMs can be extracted. As a result, it has been shown in [26] that only limited models may give a contribution to EDMs from the rainbow diagrams. In this case, the contribution from the other Feynman diagrams are thus anticipated to make the dominant contribution compared to that of the rainbow-type diagrams. The aim of the present work is therefore to revisit the previous study of Ref. [23], taking into account the diagrams that had been set aside in the belief, prior to recent developments, that only rainbow-type Feynman diagrams made the dominant contribution to charged lepton EDMs, to update our calculation and to derive a complete analytical expression for them in the case of the inverse seesaw mechanism behind active light neutrino masses and lepton mixing generation.
@@
Our calculation is performed in a non-linear Feynman-’t Hooft gauge instead of the usual gauge fixing so that the number of the contributing Feynman diagrams is considerably reduced. After having derived the complete analytical expressions for the charged lepton EDMs, we conduct a thorough numerical analysis evaluating the electron EDM, complying with all relevant experimental bounds such as neutrino oscillation data, bounds on charged lepton flavour violating processes, sterile neutrino direct searches, meson decays, sterile neutrino decays, and cosmological and astrophysical observations. For this, we have chosen the case where the most minimal ISS mechanism is embedded into the SM, that is the (2,2) ISS [29], where only two pseudo-Dirac neutrinos contribute in the loops for the charged lepton EDMs. We note that this study is general to all seesaw models leading to pseudo-Dirac neutrino pairs in the heavy spectrum. We find that the magnitude of the predicted electron EDM is $\mathcal{O}(10^{-36})~{}e\hskip 1.42271pt\mathrm{cm}$ at most, which is smaller than the previous obtained result [23]. This is because the loop function in this updated calculation is smaller than before. Nevertheless, once more sterile neutrinos are considered, there are more contribution to the electron EDM but the enhancement is only $N(N-1)/2$ times larger than in the most minimal case we considered here ($N=2$), where $N$ is the number of pairs of pseudo-Dirac neutrinos. For example, one can have a factor enhancement of 3 for $N=3$ and 6 for $N=4$. This enhancement is not enough to reach the future experimental sensitivity. In order to have an enhancement of a few orders of magnitude, one need to add a large number of sterile fermions, provided their masses are above $200$ MeV not to be in conflict with the severe constraint from Big-Bang nucleosynthesis [30]. If, in addition, one considers the ingredient fields at the origin of the LNV source, there might be further two-loop Feynman diagrams, where the heavy neutrinos as well as new neutral scalar fields run in the loops, contributing to the EDMs of charged leptons, further enhancing the phenomenological consequences of models involving sterile neutral leptons, see for instance [31].
@@ 3 Experimental constraints
Various constraints such as collider searches, meson decays, sterile neutrino decays, cosmological and astrophysical constraints are summarized in, for instance, Refs. [40, 41]. [^5] A recent study on the non-unitarity of the leptonic mixing matrix gives a preferable region for the sterile fermion mass above the electroweak scale [42]. These constraints are translated into the upper bounds for the sterile neutrino mixing matrix elements $|U_{\alpha i}|^{2}$ as a function of the sterile neutrino masses. The bounds in Ref. [40] are interpreted as a pair of nearly degenerate states for the pseudo-Dirac sterile fermions, namely $|U_{\alpha 4}|^{2}+|U_{\alpha 5}|^{2}$ and $|U_{\alpha 6}|^{2}+|U_{\alpha 7}|^{2}$ in our case.
@@ 4.2 Analytical derivation of charged lepton EDM formulae
In the (approximate) lepton number conserving case, like the models with pseudo-Dirac fermions, there are three topological diagrams relevant to the charged lepton EDMs as shown in Fig. 1, where a photon can attach to any charged particles. [^6] For lepton number violating models such as Type-I seesaw, there are additional diagrams contributing to charged lepton EDMs [22]. In fact and following Ref. [51], one can show that the contribution from the third diagram of Fig. 1, that we name “rainbow” diagram, is exactly zero by virtue of the Ward-Takahashi identity in quantum electrodynamics. Indeed, it has been recently shown that the contribution from such rainbow diagrams vanishes exactly in some cases [27, 26]. In this work, we first check that the cancellation also occurs and is exact in our model with approximate lepton number symmetry and then we proceed to calculate the contribution of the remaining diagrams. The easiest way to show this cancellation in our case could be to perform the calculation using the unitary gauge, otherwise one has to evaluate a large number of diagrams and show the exact cancellation between them. Nevertheless, and for the sake of completeness, we have checked that the exact cancellation also occurs in the non-linear Feynman-’t Hooft gauge. However, the latter calculation is tedious and time-consuming, and not worth showing in this study. The relevant rainbow Feynman diagrams are listed in Fig. 2, and the exact cancellations occur in each pair of diagrams R1–R4.
@@
The $h$ and $Z$ mediating loop contributions, are then given by
@@ 5 Numerical calculations
Fig. 6 shows the obtained CP phase factor $|J_{\alpha}^{D}|$ for the electron case ($\alpha=e$) satisfying all the relevant constraints as a function of the lightest sterile fermion mass $m_{4}~{}(\approx m_{5})$ for Normal (left) and Inverted (right) orderings. One can find that the CP phase factor for the Inverted ordering case can be about four times larger than the one in the Normal ordering case, and the maximal magnitude is $\mathcal{O}(10^{-8})$ around $m_{4}\approx 100~{}\mathrm{GeV}$. For sterile fermion masses heavier than $\mathcal{O}(1)~{}\mathrm{TeV}$, the charged lepton flavour violating processes, in particular $\mu\to e\overline{e}e$, strongly constrain the parameter space due to the non-decoupling effects [39] and the perturbative unitarity bound. This strong bound can be partially relaxed if $|\mu_{ij}|\geq 10~{}\mathrm{keV}$ is chosen. For $m_{4}\lesssim 10~{}\mathrm{GeV}$, the parameter region is also strongly limited by the beam dump experiments [40, 53], and thus we do not consider this region in our calculations.
@@ 6 Conclusions
CP-violating phases in the lepton sector are important for generating the baryon asymmetry of the Universe via leptogenesis and are linked to the mechanisms for generating small neutrino masses and their mixing as observed in neutrino oscillation experiments. In this work, we considered the charged lepton EDMs at the two-loop level as a CP-violating observable in neutrino models based on approximate lepton number symmetry and thus presenting pseudo-Dirac sterile fermions in the heavy neutrino spectrum. We found that a number of diagrams relevant to the charged lepton EDMs exactly cancel (rainbow Feynman diagrams), and we computed all contributing two-loops diagrams in the non-linear Feynman-’t Hooft gauge to derive analytic formulae for the charged lepton electric dipole moments.
@@
In addition, we numerically evaluated the predicted charged lepton EDMs taking into account the current experimental constraints such as the neutrino oscillation data, bounds on the charged lepton flavour violating processes $\ell_{\alpha}\to\ell_{\beta}\gamma$, $\alpha\neq\beta=\tau,\mu,e$ and $\mu\to e\overline{e}e$, sterile neutrino direct searches, meson decays, sterile neutrino decays, cosmological and astrophysical observations. As a result, we found that the predicted electron EDM is at most $\sim 5\times 10^{-36}~{}e\hskip 1.42271pt\mathrm{cm}$, which is too small to be explored by near future experiments. For muon and tau, the predicted EDMs are also much smaller than the experimental reaches. On the other hand, there are novel proposals aiming at measuring the electron EDM as small as $|d_{e}|<\mathcal{O}(10^{-35}-10^{-37})~{}e\hskip 1.42271pt\mathrm{cm}$ [5].
@@
This result, as is the analytical formula we derive, is generic to any fermionic seesaw based on approximate lepton number symmetry and leading to two pseudo-Dirac pairs in the heavy neutral spectrum, like, for instance, the most minimal inverse seesaw or linear seesaw mechanisms. Nevertheless, if we consider a complete model in which the seesaw mechanism is embedded, there are other two-loop contributions that will enhance these predictions. For instance, if more sterile neutrinos are considered, there are more contribution to the charged lepton EDMs but the enhancement is only $N(N-1)/2$ times larger than in the most minimal case (with $N=2$), where $N$ is the number of pseudo-Dirac pairs, meaning a factor enhancement of 3 for $N=3$ and 6 for $N=4$, which is not enough to reach the future experimental sensitivity. In order to have an enhancement of a few orders of magnitude, one needs to add a large number of sterile fermions, provided their masses are above $\sim 200$ MeV not to be in conflict with the severe constraint form Big-Bang nucleosynthesis [30]. If, in addition, one considers a more complete model with the other ingredient fields at the origin of the lepton number violating source [54, 55], there might be further two-loop Feynman diagrams, where the heavy neutrinos, as well as, new neutral and charged scalar fields run in the loops, contributing to the charged lepton EDMs and further enhancing the predicted values.
@@ Appendix A Non-linear Feynman-’t Hooft gauge
with $\nu_{\alpha}=U_{\alpha i}\nu_{i}$, the Weinberg angle $\theta_{W}$, $e=g\sin\theta_{W}$, $m_{W}=g\langle H^{0}\rangle/\sqrt{2}=m_{Z}\cos\theta_{W}$ and $\epsilon_{123}=+1$, the Lagrangian terms are given as follows